The State of Chiapas, Mexico, decided to fund a program for improving reading skills in elementary school students. The first cost is $250,000 now and an update amount of $100,000 every 6 years forever. Determine the perpetual equivalent annual cost at an interest rate of 12% per year.

The perpetual equivalent annual cost is determined to be -$___________
Answer #1

Note that the capitalized cost is given by CC = first cost + recurring cost (A/F, i%, n) / i%

CC = 250,000 + 100,000(A/F, 12%, 6) / 12%

= 250,000 + 100,000*0.123226/0.12

= 352,688.33

Now perpetual equivalent annual cost = CC * i%

= 352,688.33*12%

= $42,322.60

The answer is $42,322.60